# CRISP-SAM2 : SAM2 with Cross-Modal Interaction and Semantic Prompting for Multi-Organ Segmentation

### v1

@@ 1. Introduction
Efficient and accurate segmentation is crucial in modern medical and clinical applications, underpinning disease diagnosis, surgical planning, and outcome prediction (Ma et al., 2024; Cheng et al., 2024). Multi-organ segmentation, in particular, has been a significant area of interest due to its clinical significance and technical challenges. Traditionally, this task has been performed manually by skilled radiologists or internists, which requires meticulous case-by-case annotation and verification. This manual process is not only time-consuming and labor-intensive but also prone to errors due to the high level of expertise needed. To address these challenges, researchers have developed relatively simple and lightweight networks (Gibson et al., 2018; Wang et al., 2019; Zhou et al., 2019; Li et al., 2024a), which enable automatic segmentation but with limited accuracy. Recently, the Segment Anything Model (SAM) (Kirillov et al., 2023) has demonstrated impressive capabilities in segmenting diverse objects with the help of prompts such as points, bounding boxes, and coarse masks. Its derivative, SAM2 (Ravi et al., 2025), retains these capabilities while extending segmentation from image-level to video-level. Both SAM and SAM2 are trained on large-scale datasets, providing them with powerful zero-shot segmentation abilities in visual tasks.
@@ 2.1. Text-Guided Learning in SAMs and SAM2s
In specific tasks, SAM and SAM2 may struggle to accurately segment target masks due to a lack of domain-specific training data. A potential solution to this issue involves integrating relevant textual features using an adapter or projector to guide the segmentation process. For instance, DIT (Huang et al., 2024b), TP-DRSeg (Li et al., 2024b), and LGA (Hu et al., 2024a) enhance SAM by injecting task-related textual features after each transformer block within the image encoder. In contrast, other approaches such as AdaptiveSAM (Paranjape et al., 2024) introduces adapters in the mask decoder. Currently, the field of cross-modal interaction has seen significant advancements, particularly with the success of visual language models. Thus, the fusion of textual and visual embeddings is a logical step to create richer cross-modal features for guiding segmentation. RefSAM (Li et al., 2023b) employs multilayer perceptron (MLP) networks for this fusion, while SAP-SAM (Wang et al., 2024b) and Prompt-RIS (Shang et al., 2024) utilize cross-modal alignment and prompting modules. Similarly, EVF-SAM (Zhang et al., 2024) uses a multi-modal encoder to produce fused embeddings. To effectively bridge the gap between textual and visual modalities, models like SEEM (Zou et al., 2024) and Semantic-SAM (Li et al., 2025f) map texts, images, and prompts into a unified representation space, achieving semantic awareness. Moreover, Contrastive Language-Image Pre-Training (CLIP) (Radford et al., 2021), which excels in aligning and understanding semantics between texts and images, is employed in various segmentation tasks, such as open-vocabulary segmentation (Yuan et al., 2025; Pan et al., 2025; Wang et al., 2024a), and other related fields (Yu et al., 2024; Li et al., 2025a), offering valuable semantic insights.
@@ Acknowledgments
This work was supported by the Open Project Program of the State Key Laboratory of CAD&CG [No. A2410], Zhejiang University; National Natural Science Foundation of China[No. 61702146, 62076084, U20A20386, U22A2033].
@@ A.1. SAM and SAM2
Deep learning has made impressive progress in several fields, such as computer vision (Lu et al., 2023; Li et al., 2025c; Wu et al., 2025d; Song et al., 2025a; Zeng et al., 2024; Chen et al., 2024a; Lu et al., 2024a; Qian et al., 2024; Yao et al., 2025; Xiao et al., 2025a; Song et al., 2025b; Lu et al., 2024b; Yu et al., 2025; Gao et al., 2024; Zeng et al., 2025; Yao et al., 2025; Li et al., 2025b; Song et al., 2024), natural language processing (Bi et al., 2024, 2025; Chen et al., 2024c; Xiao et al., 2025b), automated control (Zhou et al., 2025a, b), time series analysis (Qiu et al., 2024; Wu et al., 2025b; Qiu et al., 2025c; Wu et al., 2025c; Qiu et al., 2025b; Liu et al., 2025a; Qiu et al., 2025a), multi-modal analysis (Wu et al., 2025a; Li et al., 2025e; Wu et al., 2025e; Li et al., 2025d), and multi-task learning (Xin et al., 2024a, b). However, there is still a lack of unified models in the field of computer vision, which is the reason why SAM and SAM2 are proposed.

### v2

@@ 1. Introduction
Efficient and accurate segmentation is crucial in modern medical and clinical applications, underpinning disease diagnosis, surgical planning, and outcome prediction (Ma et al., 2024; Cheng et al., 2024). Multi-organ segmentation, in particular, has been a significant area of interest due to its clinical significance and technical challenges. Traditionally, this task has been performed manually by skilled radiologists or internists, which requires meticulous case-by-case annotation and verification. This manual process is not only time-consuming and labor-intensive but also prone to errors due to the high level of expertise needed. To address these challenges, researchers have developed relatively simple and lightweight networks (Gibson et al., 2018; Wang et al., 2019; Zhou et al., 2019; Li et al., 2024d), which enable automatic segmentation but with limited accuracy. Recently, the Segment Anything Model (SAM) (Kirillov et al., 2023) has demonstrated impressive capabilities in segmenting diverse objects with the help of prompts such as points, bounding boxes, and coarse masks. Its derivative, SAM2 (Ravi et al., 2025), retains these capabilities while extending segmentation from image-level to video-level. Both SAM and SAM2 are trained on large-scale datasets, providing them with powerful zero-shot segmentation abilities in visual tasks.
@@ 2.1. Text-Guided Learning in SAMs and SAM2s
In specific tasks, SAM and SAM2 may struggle to accurately segment target masks due to a lack of domain-specific training data. A potential solution to this issue involves integrating relevant textual features using an adapter or projector to guide the segmentation process. For instance, DIT (Huang et al., 2024b), TP-DRSeg (Li et al., 2024f), and LGA (Hu et al., 2024a) enhance SAM by injecting task-related textual features after each transformer block within the image encoder. In contrast, other approaches such as AdaptiveSAM (Paranjape et al., 2024) introduces adapters in the mask decoder. Currently, the field of cross-modal interaction has seen significant advancements, particularly with the success of visual language models. Thus, the fusion of textual and visual embeddings is a logical step to create richer cross-modal features for guiding segmentation. RefSAM (Li et al., 2023b) employs multilayer perceptron (MLP) networks for this fusion, while SAP-SAM (Wang et al., 2024b) and Prompt-RIS (Shang et al., 2024) utilize cross-modal alignment and prompting modules. Similarly, EVF-SAM (Zhang et al., 2024) uses a multi-modal encoder to produce fused embeddings. To effectively bridge the gap between textual and visual modalities, models like SEEM (Zou et al., 2024) and Semantic-SAM (Li et al., 2025g) map texts, images, and prompts into a unified representation space, achieving semantic awareness. Moreover, Contrastive Language-Image Pre-Training (CLIP) (Radford et al., 2021), which excels in aligning and understanding semantics between texts and images, is employed in various segmentation tasks, such as open-vocabulary segmentation (Yuan et al., 2025; Pan et al., 2025; Wang et al., 2024a), and other related fields (Yu et al., 2024; Li et al., 2025a), offering valuable semantic insights.
@@ Acknowledgments
This work was supported by the Open Project Program of the State Key Laboratory of CAD&CG (No. A2410), Zhejiang University; National Natural Science Foundation of China (No. 61702146, 62076084, U20A20386, U22A2033); Guangdong Basic and Applied Basic Research Foundation (No. 2025A1515011617); Guangxi Key R&D Project (No. AB24010167).
@@ A.1. SAM and SAM2
Deep learning has made impressive progress in several fields, such as computer vision (Lu et al., 2023; Li et al., 2025c, f, 2024a; Wu et al., 2025d; Song et al., 2025a; Zeng et al., 2024; Chen et al., 2024a; Lu et al., 2024a; Qian et al., 2024; Yao et al., 2025; Xiao et al., 2025a; Song et al., 2025b; Li et al., 2024b; Lu et al., 2024b; Yu et al., 2025; Li et al., 2024c; Gao et al., 2024; Zeng et al., 2025; Yao and Chuah, 2024; Li et al., 2025b; Song et al., 2024), natural language processing (Bi et al., 2024, 2025; Chen et al., 2024c; Li et al., 2024e; Xiao et al., 2025b), automated control (Zhou et al., 2025a, b), time series analysis (Qiu et al., 2024; Wu et al., 2025b; Qiu et al., 2025c; Wu et al., 2025c; Qiu et al., 2025b; Liu et al., 2025a; Qiu et al., 2025a), multi-modal analysis (Wu et al., 2025a; Gu et al., 2025; Li et al., 2025e; Wu et al., 2025e; Li et al., 2025d), and multi-task learning (Xin et al., 2024a, b). However, there is still a lack of unified models in the field of computer vision, which is the reason why SAM and SAM2 are proposed.